Circular No. 7243

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA CIRCINI 1999 C. Jacques and L. Duczmal, Centro de Estudos Astronomicos de Mi-

nas Gerais, report the following CCD position obtained on Aug. 25.0529 UT with a 0.30-m Schmidt-Cassegrain telescope: α = 14^{h}23^{m}23^{s}.46, δ = −69^{o}08′45′′.3 (equinox 2000.0; uncertainty ± 0′′.4, using 39 stars from the Guide Star Catalogue 1.1). The star was at V = 7.64 (comparision star SAO 252717). G. J. Garradd, Loomberah, New South Wales, Australia, reports the following position end figures for N Cir 1999 from a CCD image taken on Aug. 25.3692 with a 0.45-m f/5.4 reflector, when the unfiltered magnitude was determined to be 7.4: 23^{s}.453, 45′′.31 (five ACT stars; rms residuals 0′′.08).

SAX J1712.6–3739 J. in ’t Zand and J. Heise, Space Research Organization Netherlands

(SRON), Utrecht; A. Bazzano and M. Cocchi, Istituto di Astrofisica Spaziale, CNR, Rome; and M. J. S. Smith, BeppoSAX Science Operation Center, Rome, and SRON, communicate: “During an observation with Wide Field Camera 2 on BeppoSAX on Aug. 24.2–25.5 UT, an x-ray transient was detected at α = 17^{h}12^{m}34^{s}, δ = −37^{o}38′.6 (equinox 2000.0; 99-percent confidence error radius 3′). The flux decreased from 32 to 20 mCrab (2– 9 keV). No x-ray bursts were detected. We designate the source SAX J1712.6–3739. The centroid is 0′.6 from the ROSAT all-sky survey source 1RXS J171237.1–373834, which had a flux of 1.6 mCrab (0.1–2.4 keV) during the survey.”

COMET 26P/GRIGG-SKJELLERUP S. M. P. McKenna-Lawlor, National University of Ireland at Maynooth,

reports the detection of a possible secondary component of comet 26P from energetic-particle (∼ 60–100 keV) observations by the EPONA instrument aboard the Giotto spacecraft on 1992 July 10. The possible secondary component was located 9 × 10^{4} km (corresponding to a delay of 40 min along the comet’s orbit) from the primary nucleus and appeared to have had a diameter 3–4 times smaller. For further details see McKenna-Lawlor and Afonin (1999, Planet. Space Sci. 47, 557).

1999 August 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green